A 5.0 mL sample of cobalt-60 used in cancer therapy contains 2.00 mCi. How many becquerels are present? (Hint: notice that you have 2 milliCuries not Curies)
Answer #1

conversion to be used:

1 Ci = 3.7*10^10 Bq

we have:

2.00 mCi = 2.00*10^-3 Ci

= 2.00*10^-3 * 3.7*10^10 Bq

= 7.40*10^7 Bq

Answer: 7.40*10^7 Bq